Matthew is 3 times as old as jenny. In 7 years, he will be twice as old as he will be then. how old is each now.
Answer by awang1996(50)

Matthew is 3 times as old as jenny. In 7 years, he will be twice as old as he will be then. how old is each now.


3j=m
m+7=2(j+7)
3j+7=2j+14
subtract 2j from both sides.
j+7=14
subtract 7 from both sides.
j=7
m=21